40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1523     Entry time: Sun Apr 26 02:13:18 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Two more successes of RF CARM handoff 
Tonight, the RF CARM hand off (mostly) succeeded twice.
But still, the IFO lost lock when I reduced the REFL_DC gain in the AO path to zero.

At the beginning of tonight's work, MICH DD hand off failed several times. This was because the the PD9 gains were set to zero.
I found that the offset script, which I called before starting the locking, fails to restore the gain values sometimes.
This happens when ezcaread fails to read the current gain. We have to be careful when running the LSCoffsets script.
ELOG V3.1.3-